# Investigating Imperceptibility of Adversarial Attacks on Tabular Data: An Empirical Analysis

### v1

@@ Abstract
Adversarial attacks are a potential threat to machine learning models, as they can cause the model to make incorrect predictions by introducing imperceptible perturbations to the input data. While extensively studied in unstructured data like images, their application to structured data like tabular data presents unique challenges due to the heterogeneity and intricate feature interdependencies of tabular data. Imperceptibility in tabular data involves preserving data integrity while potentially causing misclassification, unlike imperceptibility in images, which deals with subtle changes in pixel values. This distinction underscores the need for tailored imperceptibility criteria for tabular data. However, there is currently a lack of standardised metrics for assessing adversarial attacks specifically targeted at tabular data.
To address this gap, we derive a set of properties for evaluating the imperceptibility of adversarial attacks on tabular data. These properties are defined to capture seven perspectives of perturbed data: proximity to the original input, sparsity of alterations, deviation to datapoints in the original dataset, sensitivity of altering sensitive features, immutability of perturbation, feasibility of perturbed values and intricate feature interdepencies among tabular features.
Furthermore, this paper conducts both quantitative empirical evaluation and case-based qualitative examples analysis for seven imperceptibility properties. The evaluation reveals a trade-off between the success of adversarial attacks and imperceptibility, particularly concerning proximity, sensitivity, and deviation. Although no evaluated attacks can achieve optimal effectiveness and imperceptibility simultaneously, unbounded attacks, such as C&W and DeepFool, prove to be more promised for tabular data, emphasising their effectiveness in crafting imperceptible adversarial examples. The study also highlights the limitation of evaluated algorithms in controlling sparsity effectively. We suggest incorporating a sparsity metric in future attack algorithms to regulate the number of perturbed features. The insights gained from these empirical findings provide valuable knowledge that guides the development of imperceptible adversarial attack algorithms, proving essential for navigating the complex landscape of adversarial machine learning on tabular data.
@@ 2.2 State-of-the-art Adversarial Attacks on Tabular Data
In the context of tabular data, human eyes tend to more readily detect perturbations in feature values compared to image data. Distance-based metrics alone often cannot fully encompass all the characteristics outlined in Section 1, and additional factors must be considered. For example, the number of features perturbed (or called perturbation size) can also serve as a useful indicator for measuring imperceptibility (Ballet et al., 2019). Recent studies have also emphasised semantic imperceptibility of adversarial attacks in the context of tabular data (Mathov et al., 2022; Chernikova and Oprea, 2022). These works leverage domain knowledge to enhance the efficacy of their algorithms, strategically choosing features to perturb. Mathov et al. (2022) propose rules for immutable features and consider data types for different features. They use an embedding function incorporating all these rules to maintain the validity of perturbed examples. Similarly, Chernikova and Oprea (2022) translate domain knowledge related to network traffic into constraints, ensuring that the generated adversarial attacks do not yield in invalidated adversarial examples. These approaches showcase a subtle integration of domain expertise, shedding light on the intricacies of semantic imperceptibility of adversarial attacks in the context of in tabular data.
@@ 3.2.5 Immutability
Immutable features are fixed attributes within a dataset that are either inherently unchangeable or should not be altered due to ethical or practical considerations. Examples include personal identifiers, demographic information, and genetic data, which provide a stable foundation for analyses and ensure fairness by preventing biased alterations. Preserving the immutability of these features is crucial for maintaining data integrity and ethical compliance. Accordingly, we propose immutability as a qualitative property to evaluate whether an attack algorithm can generate adversarial examples that preserve immutable features in tabular data, which contributes to addressing the criterion of preservation of feature semantics.
@@ 3.2.6 Feasibility
We employ feasibility as a qualitative property to assess whether a perturbed feature value aligns with semantic correctness. It often requires to leverage domain knowledge, if available, or common or practical knowledge relevant to the tabular dataset for checking the practical validity of a perturbed feature value. In existing research, feasibility is introduced through the identification of feasible counterfactual examples (Poyiadzi et al., 2020) and adversarial examples (Chernikova and Oprea, 2022).
@@ 3.2.7 Feature Interdependency
Unlike image data, where pixel values are spatially arranged and often interpreted in a relatively straightforward manner by deep learning models, tabular data involves features that may have non-linear and context-specific interactions. Adversarial attacks on tabular data should navigate these intricate dependencies to produce meaningful and realistic perturbations. For instance, modifying a feature like ‘income’ in a financial dataset must be consistent with changes in related features such as ‘age’ and ‘education level’ to maintain data integrity and realism. However, simply altering ‘income’ independently of other correlated features could produce anomalies that can be easily detected.
Hence, we propose feature interdependency as a qualitative property to measures how well adversarial attack algorithms handle the intricate relationships and dependencies between features within a tabular dataset, thus addressing the criterion of preservation of feature interdependencies.
@@ 4.1 Design of Experiments
We conduct a series of experiments to evaluate the imperceptibility of adversarial attack algorithms on tabular datasets, using the properties of imperceptibility proposed in the previous section. Our experiments follow the typical machine learning pipeline, comprising data preparation, model training, and model testing. We generate adversarial examples using adversarial attack algorithms and then assess these algorithms through the following experiments:
@@
    Experiment 1: Evaluating the effectiveness of applied adversarial attack algorithms, specifically how successful these algorithms are in misclassifying a trained machine learning model.
@@
    Experiment 2: Evaluating the imperceptibility of applied adversarial attack algorithms using quantitative properties, such as sparsity, proximity, deviation, and sensitivity.
@@
    Experiment 3: Analysing case-based examples to assess the qualitative properties, such as immutability, feasibility, and feature interdependency, for the imperceptibility of applied adversarial attack algorithms.
@@
Table 1 presents the main characteristics of five datasets. All datasets are preprocessed to be suitable for binary classification. Each dataset is divided into two groups: training set (80%) used to fit the data to different machine learning models and test set (20%) to evaluate these models and generate adversarial examples. Given that the majority of distance-based adversarial attack algorithms employ the $\ell_{p}$ norm for perturbation size calculation, we utilise a one-hot encoding approach for categorical features to facilitate distance calculation. Ordinal encoding is avoided in our design, as it would introduce an unintended order within categorical features, leading to the potential distortion of distance interpretations.
@@
To the best of our knowledge, no systematic review or benchmark exists for adversarial attacks on tabular data. In this work, we review existing attack benchmarks on image-based attacks and identify those that can be adapted for tabular data. Additionally, most existing attack methods are designed for white-box settings, which means that the attacker has complete knowledge about the target predictive model thus allowing the attacker to craft effective adversarial examples to full the model. Considering these factors, our selection of attack methods is guided by the following criteria:
@@
Besides these criteria, we also consider the variety of attack methods, including both bounded attacks and unbounded attacks. Bounded attacks have an upper bound constraint $\epsilon$ on the attack budget, and this constraint limits the magnitude of allowable perturbation that the attack can make to the input data. The goal is to find an adversarial example $\bm{x}^{adv}$, which has perturbation $\bm{\delta}$ within the budget $\epsilon$, to an input $\bm{x}$ that misleads the prediction by maximising the loss function of the classifier being attacked. Conversely, unbounded attacks do not consider attack budget, and they attempt to obtain the minimal perturbation $\bm{\delta}$ between input $\bm{x}$ and adversarial example $\bm{x}^{adv}$ that misleads the prediction.
@@
We evaluate five adversarial attack algorithms on tabular data, including two bounded attacks (FGSM and PGD) and three unbounded attacks (DeepFool, LowProFool and C&W Attack). FGSM, PGD, DeepFool and C&W attack are primarily developed for image data. However, their applicability to tabuar data has been under-explored in the existing literature. Nevertheless, theoretical considerations (Goodfellow et al., 2015; Moosavi-Dezfooli et al., 2016), such as the mathematical generalisation of attack mechanisms and their algorithmic adaptability, suggest that these methods can be adapted for perturbing tabular datasets. The LowProFool algorithm is specifically tailored for tabular data, with a particular emphasis on numerical features, while neglecting categorical features in its design. As a result, this study will employ LowProFool for the evaluation of two datasets that contain only numerical data.
@@
To determine the optimal attack configurations, we perform a grid search to find the best hyperparameters and configurations for each attack method. These parameters aim to optimise the performance and proximity of the respective attack algorithms under investigation. More specifically, for two bounded attacks (FGSM and PGD), the attack budget $\epsilon$ is set to 0.3, and the perturbation size is measured using $\ell_{\infty}$ norm. PGD introduces an additional parameter known as the step size $\alpha$ to control the magnitude of each incremental step during the iterative generation of input perturbation for attack. In our experiments, $\alpha$ is set to 0.1.
@@
For three unbounded attacks (DeepFool, LowProFool and C&W attack), the minimal input perturbation is measured, using $\ell_{2}$ norm, across all three attacks. DeepFool is configured with a maximum iteration limit of $100$ and an overshoot parameter of $61e-$ 6, strategically chosen to strike a balance between maximising the attack’s effectiveness and computational costs. For C&W attack, the initial constant $c$ is set to 0.01 to control the perturbation magnitude, and a maximum of $10$ iterations is performed each involving $10$ search steps to efficiently find the minimal adversarial perturbation. C&W attack also has a confidence parameter, which influences the confidence level of misclassification, and it is set to 0 in our experiments to avoid any impact of this parameter to the attack effectiveness. LowProFool is configured with a maximum iteration limit of 100, and the feature importance is calculated using the Pearson correlation of numerical features (i.e., the default configuration). It also has a trade-off factor $\lambda$ in loss function used to balance misclassification and perturbation size, and $\lambda$ is set to 0.5.
@@
We use three representative machine learning classification models: linear support vector classification (LinearSVC) [^2] The kernelised Support Vector Classifier (SVC) addresses problems through quadratic programming and does not support loss backpropagation. In contrast, LinearSVC employs hinge loss as its optimisation criterion., logistic regression (LR) and multi-layer perceptron (MLP), to examine how different models (e.g., linear vs. non-linear models) affect the performance of adversarial attacks. LR is transparent and interpretable, whereas the decision-making process of LinearSVC and MLP is opaque to human. In order to identify potential adversarial examples, all chosen white-box attacks require the utilisation of loss gradients to back-propagate the loss function. Hence, even though tree-based models like XGBoost have demonstrated strong performance on tabular data (Shwartz-Ziv and Armon, 2022), they are excluded from the experimental design due to their inability to facilitate propagation.
@@
A higher attack success rate indicates that the assessed attack is more effective. An adversarial attack should achieve a success rate above a specified threshold before its imperceptibility can be evaluated. Subsequently, to evaluate the imperceptibility of adversarial attacks, we use quantitative metrics for key properties such as proximity, sparsity, deviation and sensitivity. We generate multiple adversarial examples for each combination of dataset and model, and then calculate the mean values for these four metrics. Additionally, the qualitative properties of immutability and feasibility are assessed through the analysis of case-based examples, which offer insights into practical implications. All the metrics used for the imperceptibility evaluation are detailed in Section 3.
@@ 4.2 Evaluation of Attack Effectiveness
To prepare for evaluating adversarial attacks, the initial phase involves training three distinct predictive models on five datasets. As shown in Table 2, subsequent testing confirms that each model achieves satisfactory accuracy, consistently exceeding the 70% threshold across all datasets. Comparing the models within the same dataset, they demonstrate similar performance levels. These findings validate the feasibility of applying adversarial attacks to these models, and it is viable to assess the effectiveness and imperceptibility of the same adversarial examples across the three models.
Table 2: Model accuracy, precision and recall of three predictive models (LR, LinearSVC, and MLP) on five datasets.
@@
Table 3: Results of attack success rates of five attack algorithms (FGSM, PGD, C&W, DeepFool, and LowProFool) on three predictive models (LR, MLP and LinearSVC) over five datasets (Adult, German, COMPAS, Diabetes, and Breast Cancer). Best attack success rate for each model is shown in bold. Overall, DeepFool is the most effective, and so is LowProFool but applicable to numerical dataset only. FGSM and PGD are generally effective and, in most cases, comparable to DeepFool. C&W is the least effective and perform differently on different models across different datasets.
@@
In contrast, C&W attack yields the lowest attack success rates among all. Especially, it demonstrates a consistently low success rate, falling below 30%, on the LinearSVC model across all datasets. One possible reason for this is the optimisation algorithms used in LinearSVC are specifically tailored for the hinge loss function. Unlike the other two predictive models LR and MLP, the hinge loss function lacks similar differentiability properties, rendering the C&W attack ineffective against LinearSVC.

### v3

@@ Abstract
Adversarial attacks are a potential threat to machine learning models by causing incorrect predictions through imperceptible perturbations to the input data. While these attacks have been extensively studied in unstructured data like images, applying them to structured data, such as tabular data, presents new challenges. These challenges arise from the inherent heterogeneity and complex feature interdependencies in tabular data, which differ from the characteristics of image data. To account for this distinction, it is necessary to establish tailored imperceptibility criteria specific to tabular data. However, there is currently a lack of standardised metrics for assessing the imperceptibility of adversarial attacks on tabular data.
To address this gap, we propose a set of key properties and corresponding metrics designed to comprehensively characterise imperceptible adversarial attacks on tabular data. These are: proximity to the original input, sparsity of altered features, deviation from the original data distribution, sensitivity in perturbing features with narrow distribution, immutability of certain features that should remain unchanged, feasibility of specific feature values that should not go beyond valid practical ranges, and feature interdependencies capturing complex relationships between data attributes.
We evaluate the imperceptibility of five adversarial attacks, including both bounded attacks and unbounded attacks, on tabular data using the proposed imperceptibility metrics. The results reveal a trade-off between the imperceptibility and effectiveness of these attacks. The study also identifies limitations in current attack algorithms, offering insights that can guide future research in the area. The findings gained from this empirical analysis provide valuable direction for enhancing the design of adversarial attack algorithms, thereby advancing adversarial machine learning on tabular data.
@@ 2.2 State-of-the-art Adversarial Attacks on Tabular Data
In the context of tabular data, human eyes tend to more readily detect perturbations in feature values compared to image data. Distance-based metrics alone often cannot fully encompass all the characteristics outlined in Section 1, and additional factors must be considered. For example, the number of features perturbed (or called perturbation size) can also serve as a useful indicator for measuring imperceptibility (Ballet et al., 2019). Recent studies have also emphasised semantic imperceptibility of adversarial attacks in the context of tabular data (Mathov et al., 2022; Chernikova and Oprea, 2022). These works leverage domain knowledge to enhance the efficacy of their algorithms, strategically choosing features to perturb. Mathov et al. (2022) propose rules for immutable features and consider data types for different features. They use an embedding function incorporating all these rules to maintain the validity of perturbed examples. Similarly, Chernikova and Oprea (2022) translate domain knowledge related to network traffic into constraints, ensuring that the generated adversarial attacks do not produce invalid adversarial examples. These approaches showcase a subtle integration of domain expertise, shedding light on the intricacies of semantic imperceptibility of adversarial attacks in the context of in tabular data.
@@ 3.2.5 Immutability
Immutable features are fixed attributes within a dataset that are either inherently unchangeable or should not be altered due to ethical or practical considerations. Examples include personal identifiers, demographic information (such as age or gender), and genetic data, which provide a stable foundation for analyses and ensure fairness by preventing biased alterations. For instance, in a medical dataset, a person’s date of birth is an immutable feature that should not be manipulated, as altering it could lead to incorrect conclusions or ethical concerns. Preserving the immutability of these features is crucial for maintaining data integrity and ethical compliance. Accordingly, we propose immutability as a qualitative property to evaluate whether an attack algorithm can generate adversarial examples that preserve immutable features in tabular data, which contributes to addressing the criterion of preservation of feature semantics.
@@ 3.2.6 Feasibility
We employ feasibility as a qualitative property to assess whether a perturbed feature value aligns with semantic correctness. It often requires to leverage domain knowledge, if available, or applying common or practical knowledge relevant to the tabular dataset for checking the practical validity of a perturbed feature value. For example, the age of a person should not be perturbed to an unrealistic value, such as 200 years. In existing research, feasibility is introduced through the identification of feasible counterfactual examples (Poyiadzi et al., 2020) and adversarial examples (Chernikova and Oprea, 2022).
@@ 3.2.7 Feature Interdependency
Unlike image data, where pixel values are spatially arranged and often interpreted in a relatively straightforward manner by deep learning models, tabular data involves features that may have non-linear and context-specific interactions. Adversarial attacks on tabular data should navigate these intricate dependencies to produce meaningful and realistic perturbations. For instance, modifying a feature like income in a financial dataset must be consistent with changes in related features such as age and education level to maintain data integrity and realism. However, simply altering income independently of other correlated features could produce anomalies that can be easily detected.
Hence, we propose feature interdependency as a qualitative property to measure how well adversarial attack algorithms handle the intricate relationships and dependencies between features within a tabular dataset, thus addressing the criterion of preservation of feature interdependencies.
@@ 4.1 Design of Experiments
We conduct a series of experiments to evaluate the imperceptibility of adversarial attack algorithms on tabular datasets, using the properties of imperceptibility proposed in the previous section. Our experiment design follows the typical machine learning pipeline, comprising data preparation, model training, and model testing. We generate adversarial examples using adversarial attack algorithms and then assess these algorithms through the following experiments:
@@
    Experiment 1: Evaluating the effectiveness of the applied adversarial attack algorithms, specifically how successful these algorithms are in misclassifying a trained machine learning model.
@@
    Experiment 2: Evaluating the imperceptibility of the applied adversarial attack algorithms using quantitative properties, including sparsity, proximity, deviation, and sensitivity.
@@
    Experiment 3: Analysing case-based examples to assess the qualitative properties, including immutability, feasibility, and feature interdependency, for the imperceptibility of the applied adversarial attack algorithms.
@@
Table 1 presents the main characteristics of five datasets. All datasets are preprocessed to be suitable for binary classification. Each dataset is divided into two groups: training set (80%) used to fit the data to different machine learning models and test set (20%) to evaluate these models. The test set will also be used to generate adversarial examples. For the same dataset, the number of adversarial examples remains consistent across different models and attacks. Given that the majority of distance-based adversarial attack algorithms employ the $\ell_{p}$ norm for perturbation size calculation, we utilise a one-hot encoding approach for categorical features to facilitate distance calculation. Ordinal encoding is avoided in our design, as it would introduce an unintended order within categorical features, leading to the potential distortion of distance interpretations.
@@
To the best of our knowledge, no systematic review or benchmark exists for adversarial attacks on tabular data. To select attack methods for evaluation, we review existing attack benchmarks on image-based attacks and identify those that can be adapted for tabular data. Additionally, most existing attack methods are designed for white-box settings, which means that the attacker has complete knowledge about the target predictive model thus allowing the attacker to craft effective adversarial examples to fool the model. Considering these factors, our selection of attack methods is guided by the following criteria:
@@
  3. 3.
    The selected attack methods should have all essential inputs available in the evaluation.
  4. 4.
    The selected attack methods should be reproducible using open-source code.
Due to criteria 3 and 4, We are unable to include two state-of-the-art methods for adversarial attacks on tabular data from our evaluation. Mathov et al. (2022) does not provide open-source code, making it non-reproducible and incompatible with our requirement for transparency and reproducibility. Chernikova and Oprea (2022) requires both datasets and specific constraints as inputs for the algorithm. These constraints rely on detailed background information and domain expertise, which are not available for our datasets sourced from the UCI repository. The absence of such domain knowledge makes it impractical for the evaluation.
@@
Besides these criteria, we also consider the variety of attack methods, including both bounded attacks and unbounded attacks. Bounded attacks have an upper bound constraint $\epsilon$ on the attack budget, and this constraint limits the magnitude of allowable perturbation that the attack can make to the input data. The goal is to find an adversarial example $\bm{x}^{adv}$, which has perturbation $\bm{\delta}$ within the budget $\epsilon$ to an input $\bm{x}$, that misleads the prediction of the model being attacked. Conversely, unbounded attacks do not consider attack budget, and they attempt to obtain the minimal perturbation $\bm{\delta}$ between input $\bm{x}$ and adversarial example $\bm{x}^{adv}$ that misleads the prediction of the model being attacked.
@@
We evaluate five adversarial attack algorithms on tabular data, including two bounded attacks (FGSM and PGD) and three unbounded attacks (DeepFool, LowProFool and C&W Attack). FGSM, PGD, DeepFool and C&W attacks are primarily developed for image data, whereas their applicability to tabuar data has been under-explored in the existing literature. Nevertheless, theoretical considerations (Goodfellow et al., 2015; Moosavi-Dezfooli et al., 2016), including mathematical generalisation of attack mechanisms and their algorithmic adaptability, suggest that these methods can be adapted for perturbing tabular datasets. The LowProFool algorithm is specifically tailored for tabular data, with a particular emphasis on numerical features, while neglecting categorical features in its design. As a result, this study will employ LowProFool for the evaluation of two datasets that contain only numerical data.
To determine the optimal attack configurations, we perform a grid search by systematically exploring a range of hyperparameter values for each attack method. This involves evaluating different combinations of hyperparameters to identify the settings that yield the best performance for each algorithm. Through this search process, we aim to optimise both the effectiveness and proximity of the attack methods under investigation. The best hyperparameters identified through this process are presented in Table 2.
Table 2: Summary of hyperparameters for bounded attack methods (FGSM, PGD) and unbounded attack methods (DeepFool, LowProFool, C&W)
Attack Hyperparameter Value
FGSM Attack Budget ($\epsilon$) 0.3
Norm $\ell_{\infty}$
PGD Attack Budget ($\epsilon$) 0.3
Norm $\ell_{\infty}$
Step Size ($\alpha$) 0.1
DeepFool Norm $\ell_{2}$
Max Iterations 100
Overshoot $61e-$ 6
C&W Norm $\ell_{2}$
Initial Constant ($c$) 0.01
Max Iterations 10
Search Steps 10
Confidence 0
LowProFool Norm $\ell_{2}$
Max Iterations 100
Trade-off Factor ($\lambda$) 0.5
Feature Importance Pearson Correlation
@@
More specifically, for two bounded attacks (FGSM and PGD), the attack budget $\epsilon$ is set to 0.3, and the perturbation size is measured using $\ell_{\infty}$ norm. PGD introduces an additional parameter known as the step size $\alpha$ to control the magnitude of each incremental step during the iterative generation of input perturbation for an attack. In our experiments, $\alpha$ is set to 0.1.
@@
For three unbounded attacks (DeepFool, LowProFool and C&W attack), the minimal input perturbation is measured, using $\ell_{2}$ norm, across all three attacks. DeepFool is configured with a maximum iteration limit of $100$ and an overshoot parameter of $61e-$ 6, strategically chosen to strike a balance between maximising the attack’s effectiveness and computational costs. For the C&W attack, the initial constant $c$ is set to 0.01 to control the perturbation magnitude, and a maximum of $10$ iterations is performed each involving $10$ search steps to efficiently find the minimal adversarial perturbation. The C&W attack also has a confidence parameter, which influences the confidence level of misclassification, and it is set to 0 in our experiments to avoid any impact of this parameter to the attack effectiveness. LowProFool is configured with a maximum iteration limit of 100, and the feature importance is calculated using the Pearson correlation of numerical features (i.e., the default configuration). It also has a trade-off factor $\lambda$ in loss function used to balance misclassification and perturbation size, and $\lambda$ is set to 0.5.
@@
We use three representative machine learning classification models: linear support vector classification (LinearSVC) [^2] The kernelised Support Vector Classifier (SVC) addresses problems through quadratic programming and does not support loss backpropagation. In contrast, LinearSVC employs hinge loss as its optimisation criterion., logistic regression (LR) and multi-layer perceptron (MLP), to examine how different models (e.g., linear vs. non-linear models) may affect the performance of adversarial attacks. LR is transparent and interpretable, whereas the decision-making process of LinearSVC and MLP is opaque to human. It is also worth noting that to identify potential adversarial examples, all chosen white-box attacks require the utilisation of loss gradients to back-propagate the loss function. Hence, even though tree-based models like XGBoost have demonstrated strong performance on tabular data (Shwartz-Ziv and Armon, 2022), they are excluded from the experimental design due to their inability to facilitate back-propagation.
@@
A higher attack success rate indicates that the assessed attack is more effective. An adversarial attack should achieve a success rate above a specified threshold before its imperceptibility can be evaluated. Subsequently, to evaluate the imperceptibility of adversarial attacks, we use quantitative metrics for key properties including proximity, sparsity, deviation and sensitivity. We generate multiple adversarial examples for each combination of dataset and model, and then calculate the mean values for these four metrics. Additionally, the qualitative properties, including immutability, feasibility, and feature interdependency, are assessed through the analysis of case-based examples, offering insights into the practical implications of adversarial attack imperceptibility. All the metrics used for the imperceptibility evaluation are detailed in Section 3.
@@ 4.2 Evaluation of Attack Effectiveness
To prepare for evaluating adversarial attacks, the initial phase involves training three distinct predictive models on five datasets. As shown in Table 3, subsequent testing confirms that each model achieves satisfactory accuracy, consistently exceeding the 70% threshold across all datasets. Comparing the models within the same dataset, they demonstrate a similar level of performance. These findings validate the feasibility of applying adversarial attacks to these models, and it is viable to assess the effectiveness and imperceptibility of the same adversarial examples across the three predictive models.
Table 3: Model accuracy, precision and recall of three predictive models (LR, LinearSVC and MLP) on five datasets.
@@
Table 4: Results of attack success rates of five attack algorithms (FGSM, PGD, C&W, DeepFool and LowProFool) on three predictive models over five datasets. The best attack success rate for each model is shown in bold. Overall, DeepFool is the most effective, and so is LowProFool but applicable to numerical datasets only. FGSM and PGD are generally effective and, in most cases, comparable to DeepFool. C&W is the least effective and performs variably across different models and datasets.
@@
In contrast, the C&W attack yields the lowest attack success rates among all. Especially, it demonstrates a consistently low success rate, falling below 30%, on the LinearSVC model across all datasets. One possible reason for this is the optimisation algorithms used in LinearSVC are specifically tailored for the hinge loss function. Unlike the other two predictive models LR and MLP, the hinge loss function lacks similar differentiability properties, rendering the C&W attack ineffective against LinearSVC.